Write an equation in standard form for the line described. Through (5,9), perpendicular to 8x+y=2
Answer by Cromlix(4381)

Hi there,
Equation: 8x + y = 2
Sort into y = mx + c
y = -8x + 2
Gradient = -8
Lines that are perpendicular
to each other have gradients
that multiply together to give
an answer of -1
m1 x m2 = -1
In this case;
-8 x m2 = -1
m2 = -1/-8
m2 = 1/8
This is the gradient for your
equation.
Using line formula:
y - b = m(x - a)
Using (5,9) (a,b)
y - 9 = 1/8(x - 5)
y - 9 = 1/8x - 5/8
y = 1/8x - 5/8 + 72/8 (9)
y = 1/8x + 67/8
or Multiply thro' by 8
8y = x + 67
Hope this helps:-)